OpenStudy (loser66):

find dy/dx implicitly and explicitly and then prove that they are equal. xy + ay -a+x=0 Please, help

OpenStudy (loser66):

explicitly: y(x+a) = a-x \(y =\dfrac{a-x}{a+x}\) then \(\dfrac{dy}{dx}= -(a+x)-(a-x)}{(a+x)^2}=\dfrac{-2a}{(a+x)^2}\)

OpenStudy (loser66):

Any mistake?

OpenStudy (loser66):

Implicitly: y +xy' +ay' +1=0 \(dfrac{dy}{dx}=\dfrac{-1-y}{x+a}\) oh, I think I find out my mistake. Sorry for bothering. :)

OpenStudy (sshayer):

correct